Arlington Manufacturing is contemplating replacing one of its machines with a new, more efficient machine. The old machine is being depreciated on a straight-line basis over the next 5 years. It has a book value of $200,000 and could be sold for $120,000. The replacement machine would cost $600,000 and have an expected life of 5 years, after which it could be sold for $100,000. Because of reductions in defects and material savings, the new machine would produce cash benefits of $180,000 per year before depreciation and taxes. The present value of $1 at 15% received after 5 periods at 15% is 0.49718. The present value of an annuity of $1 for 4 periods at 15% is 2.85498, and for 5 periods is 3.35216. Assuming straight-line depreciation, a 40% marginal state and federal tax rate, and a required rate of return of 15%, find the payback period and NPV.
Answer #1
a Initial Cash Flow:
Before tax sallvage value of old machine $120,000
Book value of old machine $200,000
Gain on Salvage=200000-120000= $80,000
Tax on gain =80000*40% $32,000
A After tax salvage value of old machine $88,000 (120000-32000)
B Cost of replacement machine ($600,000)
C=A+B Initial Cash Flow: ($512,000)
DEPRECIATION OF OLD MACHINE
Year from today 1 2 3 4 5
D1 Annual Depreciation expense=200000/5= $40,000 $40,000 $40,000 $40,000 $40,000
DEPRECIATION OF NEW MACHINE
Year(from today) 1 2 3 4 5
D2 Annual Depreciation expense=(600000-100000)/5 $100,000 $100,000 $100,000 $100,000 $100,000
D=D2-D1 Change in annual depreciation expense $60,000 $60,000 $60,000 $60,000 $60,000
N Year From Today 1 2 3 4 5
E=D*40% Depreciation tax shield $24,000 $24,000 $24,000 $24,000 $24,000
F Before tax saving in operating cost $180,000 $180,000 $180,000 $180,000 $180,000
G=F*(1-0.4) After tax increase in operating cash flow $108,000 $108,000 $108,000 $108,000 $108,000
H Salvage value $100,000
I=E+G+H Incremental Net Cash Flow $132,000 $132,000 $132,000 $132,000 $232,000
Cumulative cash flow
J Present Value of Year1-4 Cash Flows $376,857 (2.85498*132000)
K Present Value of Year 5 Cash Flows $115,346 (0.49718*232000)
L=J+K Present Value of Cash Inflows $492,203
I Sumof Present Value of Cash outflow -$512,000
NPV=L+I Net Present value ($19,797)
NPV is Negative
No, the machine should NOT be purchased
Payback Period=512000/132000=                 3.88 Years